I've been trying to solve the following problem for two days now....and only getting frustrated. Solve: 5 minus the (square root of x squared minus 4x plus 32) = x plus 5 NOTE: the problem asks for the square root of x squared minus 4x plus 32 and not just x squared...
Many Thanks,
Cindy

Answer by checkley71(8403)

5-SQRT(X^2-4X+32)=X+5 COMBINING THE 5 S WE GET
-SQRT(X^2-4X+32)=X NOW SQUARE BOTH SIDES WE GET
-(X^2-4X+32)=X^2
-X^2+4X-32-X^2=0 COMBINE X TERMS
-2X^2+4X-32=0 DIVIDE BY 2
X^2-2X+16=0 FACTOR USING THE QUADRATIC EQUATION
X=(-B+-SQRT[B^2-4AC])/2A
X=(2+-SQRT(-2*-2)-(4*1*16))/2*1
X=(2+-SQRT4-64)/2
X=(2+-SQRT-60)/2
X=(2+-7.746i)/2
X=(2+7.746i)/2
X=1+3.87i SOLUTION
X=(2-7.746i)/2
X=1-3.87i SOLUTION